LINE SHAPE PARAMETERS FOR CO_{2} TRANSITIONS: ACCURATE PREDICTIONS FROM COMPLEX ROBERT-BONAMY CALCULATIONS

\underline{JULIEN LAMOUROUX} AND ROBERT R. GAMACHE, Department of Environmental, Earth, and Atmospheric Sciences, University of Massachusetts Lowell, Lowell, MA 01854, USA.

A model for the prediction of the vibrational dependence of CO_{2} half-widths and line shifts for several broadeners, based on a modification of the model proposed by Gamache and Hartmann^{a}, is presented. This model allows the half-widths and line shifts for a ro-vibrational transition to be expressed in terms of the number of vibrational quanta exchanged in the transition raised to a power p and a reference ro-vibrational transition. Complex Robert-Bonamy calculations^{b} were made for 24 bands for lower rotational quantum numbers J′′ from 0 to 160 for N_{2}-, O_{2}-, air-, and self-collisions with CO_{2}. In the model a Quantum Coordinate is defined by (c_{1}Δν_{1} + c_{2}Δν_{2} + c_{3}Δν_{3})^{p} where a linear least-squares fit to the data by the model expression is made. The model allows the determination of the slope and intercept as a function of rotational transition, broadening gas, and temperature. From these fit data, the half-width, line shift, and the temperature dependence of the half-width can be estimated for any ro-vibrational transition, allowing spectroscopic CO_{2} databases to have complete information for the line shape parameters.

aR. R. Gamache, J.-M. Hartmann, J. Quant. Spectrosc. Radiat. Transfer. 83 (2004), 119.

bR. R. Gamache, J. Lamouroux, J. Quant. Spectrosc. Radiat. Transfer. 117 (2013), 93.